OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. Answer y = -3x + 3 y - 6 = -3(x + 1) y - 6 = -3x - 3 3x + y = 3

OpenStudy (anonymous):

Point Slope form is (y - y1) = m(x - x1) so we get y - 6 = -3 (x + 1) as the correct answer